Use the following information to answer questions 25-34 We all know that dogs have an amazing sense of smell. They are used tApplying the 3-s process 28. (5 points) Statistic. What is the statistic? 29. (5 points) Simulate. In your own words, explainUse the paragraph to answer 25-34 (all parts of one question)
Answer #1

25. Ho : Marine is able to indentify patients with cancer

H1 : Marine is not able to identify patients with cancer or is just guessing

26. Here, for a sample of 33 units, she was correct 30 times. Here we can compute the probability of identifying a cancer patient correctly , p = 30/33 = 0.91 and in ideal scenario we can assume population P = 1.

27. The variable is p = 0.91 and q = 0.09 , n = 33 which will follow bernaulli dist with mean p and variance pq/n.

28. The statistic will bez = \frac{p-P}{\sqrt{PQ/n}} = -1.8065 and can be compared with critical value or p-value with level of significance.